Three students approved in an examination with the following probability to pass the exam 9/10,8/10,7/10.what is the probability that atleast one of them will pass the exam
Answer by ikleyn(52810)

.
Three students approved in an examination with the following probability
to pass the exam 9/10,8/10,7/10.
what is the probability that atleast one of them will pass the exam
~~~~~~~~~~~~~~~~

The probability that no one will pass the exam is

    P = (1-0.9)*(1-0.8)*(1-0.7) = 0.1*0.2*0.3 = 0.006.


The probability that at least one will pass the exam is the complement to it

    P' = 1 - P = 1 - 0.006 = 0.994.     ANSWER

Solved.